For what values of h the vectors


⟶              ⟶                 ⟶  

u1 = [1, -3, -2]             u2 = [-1, 9, -6]              u3 = [5, -7, h]   

     

are linearly independent? (Show all working)


Note: The three vectors are supposed to be in a 3x1 matrix(3 rows and 1 column)



1
Expert's answer
2022-03-24T19:15:06-0400

"\\left| \\begin{matrix}\t1&\t\t-1&\t\t5\\\\\t-3&\t\t9&\t\t-7\\\\\t-2&\t\t-6&\t\th\\\\\\end{matrix} \\right|=\\left( 9h-42 \\right) +\\left( -3h-14 \\right) +5\\left( 18+18 \\right) =\\\\=6h+123\\\\6h+123\\ne 0\\Rightarrow h\\ne -20.5"